OpenStudy (anonymous):

Find the value of : \[\frac{1}{2 \times 3} + \frac{1}{3 \times 4} + \frac{1}{4 \times 5} + \frac{1}{5 \times 6} + ............+ \frac{1}{9 \times 10} \] Explain your answer....

OpenStudy (anonymous):

what is the '..............'?

OpenStudy (anonymous):

It means the sum goes to 1/(9*10....

OpenStudy (anonymous):

The terms in place of ............................. are: \[\frac{1}{6 \times 7} + \frac{1}{7 \times 8} + \frac{1}{8 \times 9}\]

OpenStudy (ash2326):

We have \[\frac{1}{2 \times 3} + \frac{1}{3 \times 4} + \frac{1}{4 \times 5} + \frac{1}{5 \times 6} + ............+ \frac{1}{9 \times 10}\] first term can be written as \[\frac{1}{2\times 3}=\frac 12 -\frac 13\] Using this we'll rewrite the termsas \[\frac{1}{2}-\frac 13+\frac 13-\frac 14+\frac 14-\frac 15+\frac 15-\frac 1 6+\frac 16-\frac 17+\frac 17-\frac 18+\frac 18-\frac 19+\frac19-\frac 1{10}\] Now we'll get finally \[\frac 12-\frac {1}{10}\] I think you can solve now!

OpenStudy (anonymous):

Is 7429 is a factor of 20160??

OpenStudy (anonymous):

2/5.. Yeah thanks @ash2326 a lot..

OpenStudy (ash2326):

Did you understand @waterineyes ?

OpenStudy (anonymous):

Ya I did understand.. \[\frac{1}{5 \times 6} = \frac{1}{5} - \frac{1}{6} = \frac{6-5}{30} = \frac{1}{ 5 \times 6}\]

OpenStudy (anonymous):

So every term in between cancels out each other and we are left with first and the last term...

OpenStudy (ash2326):

Yeah \[\frac{1}{2}-\cancel{\frac 13}+\cancel{\frac 13}-\frac 14+\frac 14-\frac 15+\frac 15-\frac 1 6+\frac 16-\frac 17+\frac 17-\frac 18+\frac 18-\frac 19+\frac19-\frac 1{10}\]

OpenStudy (anonymous):

Yeah I got it ashutosh.. Thanks...

OpenStudy (ash2326):

Welcome:D